OpenStudy (rissa):

cos^2 theta (1+tan^2 theta)

OpenStudy (rissa):

\[\cos ^{2}\theta (1+\tan ^{2}\theta)\]

OpenStudy (zzr0ck3r):

Hint: \(\tan(x)=\frac{\sin(x)}{\cos(x)}\) \(a(b+c)=ab+ac\) \(a\frac{a}{b}=b\) \(\cos^2(x)+\sin^2(x)=1\)

OpenStudy (rissa):

i got the answer 1

OpenStudy (rissa):

am i right ?

OpenStudy (zzr0ck3r):

yep

OpenStudy (zzr0ck3r):

I am not sure how strict your teacher is, but \(1\) as a function does not have a restricted domain, where the function on the left side does, so you may need to say something like \(..=1\) as long as \(\theta \ne \) something that makes cosine \(0\).

OpenStudy (sshayer):

\[\cos^2 \theta \left( 1+\tan ^2\theta \right)=\cos ^2\theta \sec ^2\theta=\frac{ \cos ^2\theta }{ \cos ^2\theta }=1\] if\[\cos \theta \neq 0\]

OpenStudy (zzr0ck3r):

This is another way \(\uparrow\) Using the fact that if we divide \(\cos^2(x)+\sin^2(x)=1\) on both sides by \(\cos^2(x)\) we get \(1+\tan^2(x)=\sec^2(x)\)